OpenStudy (anonymous):

suppose a rover landed on Jupiter and could launch a rock into the air. assume the function h(t)=-13^2+50t+2 models the rocks height above the surface,h, in meters after t seconds from release. what is the height of the rock at the time of release?

OpenStudy (anonymous):

Set t=0 and solve.

OpenStudy (luigi0210):

It would be 2 meters, if you set it equal to 0..

OpenStudy (anonymous):

okay thanks. the next part of the question is, when will the rock reach its maximum height ?

OpenStudy (anonymous):

Assuming, of course, that the 13^2 is supposed to mean 13 t^2.

OpenStudy (anonymous):

yes sorry 13t^2

OpenStudy (anonymous):

Maximum height is most easily found by taking the derivative and setting it equal to zero. Solve for t and you have your time, and reinsert that t in the original equation for the height.

OpenStudy (anonymous):

am i using the quadratic equation by setting t to 0?

OpenStudy (luigi0210):

You could graph it too, if you haven't learned derivatives

OpenStudy (anonymous):

No, that'd get you the time when h=0. The derivative of quadratics at^2+bt+c is 2at+b, but I think there was an elaborate equation to find h without that...

OpenStudy (agent0smith):

There's no need for calculus or graphs. Just find the vertex of the parabola from \[\Large x= -\frac{ b }{ 2a }\] (though that ^ is derived with calculus) Then plug x back into the h equation.

OpenStudy (agent0smith):

Since it's t, \[\Large t= -\frac{ b }{ 2a }\] a and b are from \[\large ax^2 +bx +c\]

OpenStudy (anonymous):

im getting a tad confused lol, i know there is a set of equations to solve this particular problem i just cant seem to remember them, there are 5 parts to this question : what is the height of the rock at the time of release, how high is the rock after one second, when will the rock reach its maximum height, what is the maximum height and when will the rock hit the surface. so you're saying that to find the height at time of release i should use the equation -b/2a?

OpenStudy (agent0smith):

when will the rock reach its maximum height when t=-b/2a

OpenStudy (anonymous):

Height at time of release is just H(0). H(1) gives height at one second. Max height is given by t=-b/2a. H(t) with that t gives the height itself. The rock hits the surface, assuming the surface is at height 0, when H(t) =-13t^2+50t+2 0 (use a quadratic here, and remember that negative times tell you when the rock would have been launched if it were to be in the same place with the same velocity as this rock but are otherwise nonsensical)

OpenStudy (anonymous):

okay thank you ! i'll go try those out now.

OpenStudy (anonymous):

hi again lol, so i got 3.88 seconds for when the rock will hit the planets surface, does that make sense to you all ? if so i did it right

OpenStudy (agent0smith):

Looks about right. You can check your values from a graph... https://www.google.com/search?q=-13t%5E2%2B50t%2B2&aq=f&oq=-13t%5E2%2B50t%2B2&aqs=chrome.0.57.1729j0&sourceid=chrome&ie=UTF-8